# BECS: A Privacy-Preserving Computing Sharing Mechanism for 6G Computing Power Network

### v2

@@ V-C Pseudonym Generation
When $D_{d}$ participates in computing sharing, it must first send its public key $dg$ and identity certificate $Cert_{d}$ to the nearest $B_{b}$. After $B_{b}$ verifies the legitimacy of $D_{d}$’s identity, it generates a pseudonym for $D_{d}$, as detailed in Algorithm 2. Based on the Schnorr protocol [36], $D_{d}$ proves possession of the private key $d$ to $B_{b}$ without revealing it, by utilizing the public parameter $g$ and random numbers. $B_{b}$ verifies the legitimacy of $D_{d}$’s identity by confirming the correctness of $D_{d}$’s proof and response. Utilizing non-interactive zero-knowledge protocol, $D_{d}$ acquires a pseudonym $(x,y)$ for communication from $B_{b}$ while preserving privacy.
@@
0:  $D_{d}$’s private key $d$, public key $dg$.
0:  $D_{d}$’s pseudonym $(x,y)$.
1:  $D_{d}\longrightarrow B_{b}$: Sends $(\tilde{x}=g,\tilde{y}=dg)$.
2:  $B_{b}\longrightarrow D_{d}$: Chooses $\gamma\in Z_{q}^{*}$, sends $x=\gamma\tilde{x}$.
3:  $D_{d}\longrightarrow B_{b}$: Calculates $y=dx$, then chooses $\delta\in Z_{q}^{*}$, and calculates commitment $K=\delta x$, challenge $\epsilon=\mathcal{H}(x\parallel y\parallel K)$, response $M=\delta+\epsilon d\mathrm{~{}mod~{}}q$, sends $y,K$, and $M$.
4:  $B_{b}$: Calculates $\epsilon^{\prime}=\mathcal{H}(x\parallel y\parallel K)$ to validates $Mx\stackrel{{\scriptstyle?}}{{=}}K+\epsilon^{\prime}y$.
5:  if the verification is correct then
@@
0:  $D_{d}$’s pseudonym $(x,y)$; $B_{b}$’s private key $b$, public key $h=bg$.
0:  $D_{d}$’s pseudonym certificate $\varpi$.
1:  $B_{b}\longrightarrow D_{d}$: Chooses $\phi\in Z_{q}^{*}$, and calculates commitment $O=\phi g$, challenge $\zeta=\mathcal{H}(x\parallel y\parallel O)$, response $P=\phi+\zeta b\mathrm{~{}mod~{}}q$, sends $O$ and $P$.
2:  $D_{d}$: Calculates $\zeta^{\prime}=\mathcal{H}(x\parallel y\parallel O)$ to verifies $Pg\stackrel{{\scriptstyle?}}{{=}}O+\zeta^{\prime}h$, if verification succeeds, then store the certificate $\varpi=(O,P)$.
@@ V-D Pseudonymous Computing Trading
Upon entering the system, CRP synchronizes its computing device information and status within blockchain. CRR sends a computing request, and NSGA-III-KDR is utilized to match the resource-sharing parties. CRP ($D_{2}$) accepts a task offloading request from CRR ($D_{1}$). Initially, both parties must verify the pseudonym and certificate according to Algorithm 4. $D_{1}$ uses transcript $T_{1}$ to prove the authenticity of the pseudonym to $D_{2}$ and uses the certificate $\varpi_{1}$ to prove the legitimacy of the certificate to $D_{2}$. After both verifications succeed, $D_{2}$ confirms $D_{1}$’s identity, and similarly, $D_{1}$ confirms $D_{2}$’s identity. Once confirming their identities, CRR offloads its computing tasks to CRP, which then updates its resource status and delivers the computing results to CRR within the deadline. Following the payment of corresponding fees by CRR, both parties finalize the trading.
@@
0:  $D_{1}$’s pseudonym $(x_{1},y_{1})$, transcript $T_{1}=(K_{1},M_{1})$, and certificates $\varpi_{1}=(O_{1},P_{1})$ are issued by $B_{1}$; $B_{1}$’s public keys $h_{1}=b_{1}g$.
0:  $D_{2}$ accepts or rejects $D_{1}$’s identity.
1:  $D_{1}\longrightarrow D_{2}$: Sends transcript $T_{1}=(K_{1},M_{1})$ and certificates $\varpi_{1}=(O_{1},P_{1})$ under the pseudonym $(x_{1},y_{1})$.
2:  $D_{2}$: Calculates $\epsilon_{1}^{\prime}=\mathcal{H}(x_{1}\parallel y_{1}\parallel K_{1})$ to verifies the authenticity of pseudonym: $M_{1}x_{1}\stackrel{{\scriptstyle?}}{{=}}K_{1}+\epsilon_{1}^{\prime}y_{1}$. Then Calculates $\zeta_{1}^{\prime}=\mathcal{H}(x_{1}\parallel y_{1}\parallel O_{1})$ to verifies the legitimacy of pseudonym and certificate: $P_{1}g\stackrel{{\scriptstyle?}}{{=}}O_{1}+\zeta_{1}^{\prime}h_{1}$.
3:  If all verifications are correct, $D_{2}$ accepts $D_{1}$’s identity. Else, $D_{2}$ rejects $D_{1}$’s identity.
@@
V-E Block Generation and Pseudonymous Update
After both parties complete the trading in computing sharing, the 6GBS near CRR packages all related information into a new block and adds it to the blockchain. Subsequently, CRP must update its computing resource status. The trading and resource update information of this computing sharing will be broadcast and synchronized throughout the entire blockchain system.
Before re-engaging in computing sharing, CRP and CRR need to request new pseudonyms and certificates from the nearby 6GBS to ensure the user’s identity remains unlinkable. Similarly, the 6GBS verifies the legitimacy of the CRP/CRR identities through Algorithm 4 and then issues CRP/CRR new pseudonyms and certificates.
@@ VI Security Analysis and Discussion
VI-A System Security
The system’s security is crucial for the anonymous sharing and trading of computing resources. In BECS, the RCC of 6GBS generates all security parameters, which are then securely synchronized and stored on the permissioned blockchain. Additionally, all 6GBSs maintain the permissioned blockchain, and their security strongly supports the blockchain. Thus, resource status updates, pseudonymous trading, trading records, and block generation within the system are secure. Moreover, as all computing trading are blockchain-recorded, 6GBSs can ensure fairness in computing sharing through audits. Furthermore, user authentication is mandatory for system access, and user identity hashes stored at the 6GBS ensure privacy even if data leaks occur, also safeguarding against whitewashing attacks. Finally, updatable pseudonyms prevent the leakage of users’ public and private keys, ensuring that their information remains unlinked and untraceable, thus securing their privacy. In the proposed service model, privacy entropy is also considered, providing comprehensive protection for user privacy.
VI-B Pseudonymous Security
The security of the proposed pseudonym scheme is based on the discrete logarithm problem and the Schnorr protocol. During system initialization, the security of key generation depends on the assumption of the discrete logarithm on elliptic curves, and its security is notoriously guaranteed. Additionally, the Schnorr protocol, being simulatable due to its zero-knowledge property [36], ensures that the pseudonym generation is also simulatable, with interacting parties not gaining any information beyond validating the pseudonyms and certificates, so the generation and verification of pseudonyms and certificates is also secure. Moreover, since attackers cannot differentiate between real and simulated interactions, they are unable to acquire any useful information. To analyze security performance from various perspectives, the detailed discussion as follows.
1) Uniqueness: The proposed pseudonym adopts the form $(\gamma g,\gamma dg)$, whereas the user’s public/private key pairs take the form $(d,dg)$. The pseudonym generation can be seen as a knowledge extractor of the Schnorr protocol that produces the private key $d$. Consequently, each user’s pseudonym corresponds to a unique extractable public/private key pairs, ensuring that every authenticated pseudonym uniquely identifies a user.
2) Unforgeability: The user’s pseudonym certificate incorporates the private key of 6GBS, which remain concealed from users due to the zero-knowledge protocol, thereby precluding any possibility of forging the pseudonym certificate. Additionally, user identity verification employs a zero-knowledge protocol based on the 6GBS’s private key. Consequently, without knowing the 6GBS’s private key, users cannot authenticate successfully with a forged pseudonym and certificate.
3) Unlinkability: The mutual verification process between users reveals no information beyond the pseudonym and certificate. The pseudonym is a bit commitment to the user’s private key. Given the robust security of bit commitments, attackers cannot distinguish the identities of different users by observing their pseudonyms. Furthermore, the public keys of both verification parties are not leaked, and no link exists between different pseudonyms of the same user.
4) Traceability: Users request or update their pseudonyms at the 6GBS using their public keys, allowing the 6GBSs to trace all of a user’s pseudonyms and related trading via its public key. As pseudonyms are updated before each computing sharing and remain unlinkable, neither attackers nor other users can trace a user’s pseudonyms or related trading.
VI-C Certificate Security
The certificate’s security is based on key pairs of the 6GBS. Certificate issuance and verification, conducted through non-interactive zero-knowledge protocols, ensuring the confidentiality of the 6GBS’s key pairs. The private key of the granting 6GBS binds the pseudonym, thereby legitimizing the certificate. Additionally, during certificate issuance, random numbers further improve anonymity, safeguarding against linking of the user’s identity.
VI-D Computational Complexity of NSGA-III-KDR
@@
In BECS, NSGA-III-KDR uses Non-KDR-dominated sort instead of Non-Dominated Sort from NSGA-III to enhance solution diversity, while maintaining the same computational complexity. Namely, with $L$ optimization objectives and a population size of $X$, NSGA-III-KDR exhibits a computational complexity of $\mathcal{O}(LX^{2})$. Specifically, calculating the kernel distance for each solution and determining the angle between any two solutions each incur a complexity of $\mathcal{O}(L)$. Each solution’s distance to the ideal point is calculated individually, contributing to a complexity of $\mathcal{O}(LX)$. For dominance judgment, each pair of solutions undergoes one angle calculation and comparison, leading to an overall complexity of $\mathcal{O}(LX^{2})$. Thus, the computational complexity of each iteration can be approximated as $\mathcal{O}(LX^{2})$. Furthermore, while maintaining the same complexity, NSGA-III-KDR supports more optimization objectives compared to NSGA-II-based scheme [11], providing a more comprehensive computing allocation strategy. Compared to NSGA-III-based scheme [12], NSGA-III-KDR offers more diverse solutions, improving the probability of superior computing allocation strategies.
@@ VII Simulation Results and Analysis
In this section, we initially compare the performance of the proposed NSGA-III-KDR with NSGA-III and MOEA/D applied in computing allocation, as well as the NSGA-II-SDR that inspired us. Subsequently, we test the performance of four algorithms in optimizing the proposed computing allocation MOOP. Finally, we simulate the performance of the proposed pseudonym scheme. The configurations of critical parameters are detailed in Table II.
TABLE II: Key Parameters
@@
CPU frequency [37] $\mathbb{U}$: 0.6 GHzsimilar-$\sim$10 GHz; $\mathbb{E}$: 10 GHzsimilar-$\sim$1 THz; $\mathbb{C}$: $>$ 1THz
@@ VII-A Simulation of Proposed NSGA-III-KDR
In this part, we compare the proposed NSGA-III-KDR with state-of-the-art evolutionary algorithms NSGA-III [11], MOEA/D [12], and NSGA-II-SDR [13], utilizing the PlatEMO platform [39]. The widely used SDTLZ [31], MaF [40], and SMOP [41] test suites are employed as the benchmarks. Meanwhile, IGD [42] and PD [43] are selected as performance evaluation metrics. IGD comprehensively measures the convergence and diversity of the algorithms, where a smaller IGD value indicates better performance. PD primarily reflects the diversity of the algorithms, with a larger PD value indicating greater population diversity. The crossover probability is set to 1, the mutation probability is set to 1/D, and their distribution indicator is set to 20, where D represents the length of the decision variable. All three algorithms are executed 30 times on different test problems, and the average values are taken. The performance of the three algorithms is compared under different numbers of objectives and benchmarks, as shown in Table III, where “$+$”, “--”, and “$=$” indicate that the result is significantly better, significantly worse, and statistically similar to that obtained by NSGA-III-KDR, respectively.
It can be concluded from the experimental results that NSGA-III-KDR has the strongest competitiveness, achieving the best IGD and PD numbers of 10 and 15, respectively. This demonstrates that NSGA-III-KDR has better performance compared to the other two algorithms, especially in terms of population diversity, offering a richer set of solutions for computing allocation.
TABLE III: IGD and PD value of NSGA-III, MOEA/D, NSGA-II-SDR, and NSGA-III-KDR on SDTLZ1, SDTLZ2, MaF1, MaF2, SMOP1, and SMOP2 with 5, 10, and 15 objectives. The best result in each row is highlighted.

### v3

@@ V-C Pseudonym Generation
When $D_{d}$ participates in computing resource sharing, it must first send its public key $dg$ and identity certificate $Cert_{d}$ to the nearest $B_{b}$. After $B_{b}$ verifies the legitimacy of $D_{d}$’s identity, it generates a pseudonym for $D_{d}$, as detailed in Algorithm 2. Based on the Schnorr protocol [31], $D_{d}$ proves possession of the private key $d$ to $B_{b}$ without revealing it, by utilizing the public parameter $g$ and random numbers. $B_{b}$ verifies the legitimacy of $D_{d}$’s identity by confirming the correctness of $D_{d}$’s proof and response. Utilizing non-interactive zero-knowledge protocol, $D_{d}$ acquires a pseudonym $(x,y)$ for communication from $B_{b}$ while preserving privacy.
@@
[…]
5:if verification succeeds then
@@
[…]
@@ V-D Pseudonymous Computing Trading
The pseudonymous computing trading in BECS is structured as a hybrid on-chain/off-chain protocol designed to ensure both real-time responsiveness and verifiable trust. This architecture decouples rapid off-chain execution from asynchronous on-chain settlement, thereby mitigating the inherent latency caused by blockchain consensus. While a CRP initially registers its identity and core capabilities on-chain to establish a foundational trust anchor, its dynamic operational status is maintained off-chain. When a CRR ($D_{1}$) initiates a request, the NSGA-III-KDR matches it with an optimal CRP ($D_{2}$). The algorithm operates on a near real-time view of resource availability, maintained by the 6GBS and continuously updated based on both the canonical on-chain state and transient off-chain “soft-state” messages. The trading process consists of two primary phases:
1) Off-Chain Execution: First, both parties verify their pseudonyms and certificates according to Algorithm 4 to establish mutual trust through a secure off-chain channel. Subsequently, they negotiate and finalize an off-chain agreement by cryptographically signing the terms of service. Importantly, upon reaching this agreement, $D_{2}$ immediately broadcasts a signed lightweight status message to the nearest 6GBS. This message serves as an off-chain “soft-state” update, temporarily marking the resource as “in use.” This rapid propagation ensures that $D_{2}$ is instantly excluded from the pool of available resources in any concurrent NSGA-III-KDR matching process, thereby preventing double booking. To ensure system resilience and prevent resource deadlocks, this soft state is linked to a predefined timeout, and the resource is automatically released if an on-chain settlement transaction is not detected within this period. This mechanism allows the CRR to offload computing tasks for immediate execution without waiting for blockchain confirmation.
2) On-chain Settlement: Second, this phase begins only after the CRP completes the task and delivers the results to the CRR. Upon successful verification of the outcome, a formal transaction is generated. This transaction includes the essential details of the trade, such as digital signatures from both parties, hashed references to the off-chain agreement, and proof of task completion. The transaction is subsequently submitted to the blockchain. The successful inclusion of this transaction in a block finalizes the trade, generating an immutable and auditable record while formally updating the canonical state of the CRP’s resources on the ledger.
V-E Block Generation and Pseudonymous Update
Following the transaction submission, the 6GBSs serve as trusted consensus nodes responsible for validating transactions and maintaining the ledger. The selection of the 6GBS responsible for block generation is determined by the Proof of Trust and Adjustment (PoTA) consensus mechanism, a lightweight and efficient protocol that we introduced in our previous work [17]. Specifically, the PoTA protocol selects a bookkeeping node based on a dynamically computed score. This score comprehensively evaluates both a node’s global trust, derived from historical interaction ratings, and its relevant resource service attributes. This mechanism maintains a balance between security and efficiency, making it a practical and effective consensus solution for dynamic 6G CPN environments. The 6GBS with the highest score is authorized to aggregate all relevant information, such as transaction details, digital signatures, and resource status updates, into a new block. This block is then broadcast for validation and subsequently appended to the blockchain.
Subsequently, the CRP’s computing resource status is formally updated and finalized within the new block, serving as the canonical on-chain record of the state change that was already broadcast off-chain as a soft state. This two-step process ensures that the NSGA-III-KDR algorithm operates on the most up-to-date information from off-chain soft-state messages, while the blockchain maintains the ultimate ground truth for consistency and security. Before re-engaging in computing resource sharing, the CRP and CRR must request new pseudonyms and certificates from the nearest 6GBS to ensure their identities remain unlinkable. Similarly, the 6GBS verifies the legitimacy of the CRP and CRR identities using Algorithm 4 and subsequently issues them new pseudonyms and certificates.
@@ V-E Block Generation and Pseudonymous Update
[…]
3: If all verifications are successful, $D_{2}$ accepts $D_{1}$’s identity. Else, $D_{2}$ rejects $D_{1}$’s identity.
@@ VI Security Analysis and Discussion
VI-A Threat Model and Adversary Assumptions
To facilitate formal analysis, a comprehensive threat model is adopted, encompassing two primary types of adversaries:
  • External Adversary ($\mathcal{A}$): An entity that is not a legitimate participant in the 6G CPN. Its capabilities are limited to eavesdropping on communication channels, with the objective of compromising user privacy via traffic analysis or forging messages to obtain unauthorized access.
  • Malicious Internal Node ($\mathcal{B}$): A registered participant within the BECS system (e.g., a compromised CRP or CRR) may comply with protocol specifications but still attempt to exploit its privileges to deanonymize transaction partners, repudiate actions, or disrupt system operations by injecting false information.
  • Statistical Inference Adversary ($\mathcal{C}$): A curious-but-honest internal participant (e.g., an MEC server) that correctly processes tasks while passively analyzing historical offloading metadata [64]. Its goal is to compromise user privacy by inferring wireless channel states from observed offloading patterns, thereby revealing the user’s physical location and behavioral characteristics.
The security analysis relies on standard cryptographic assumptions, including the computational hardness of the discrete logarithm problem and the collision resistance of the hash function $\mathcal{H}$.
VI-B System Security
The permissioned blockchain architecture establishes a foundational layer of defense against both external and internal adversaries.
1) Access Control and Authentication: This mechanism constitutes the first line of defense against the external adversary $\mathcal{A}$. By mandating that all devices register through the RCC and obtain verifiable credentials, the system effectively prevents unauthorized entities from joining the network, thereby mitigating risks such as impersonation and unauthorized access to resources.
2) Integrity and Non-repudiation: Blockchain immutability serves as a direct countermeasure to threats posed by the malicious internal node $\mathcal{B}$. It ensures that malicious participants cannot tamper with transaction records to alter agreed-upon terms or repudiate their actions. All activities are irreversibly recorded, providing a verifiable audit trail that enforces accountability.
3) Resilience and Availability: The blockchain’s distributed architecture is inherently resilient to targeted attacks by a powerful adversary $\mathcal{B}$ capable of compromising one or several 6GBSs. By eliminating single points of failure, the system ensures service continuity and preserves data integrity even in the presence of partial compromise.
4) Resistance to Statistical Inference Attacks: The privacy entropy maximization objective in the proposed MOOP provides a statistical defense against inference attacks launched by the malicious internal node $\mathcal{B}$. By promoting diverse and unpredictable task offloading behaviors, this mechanism obfuscates user behavioral patterns, thereby complementing the direct identity unlinkability provided by the pseudonym scheme.
5) Balancing Transparency and Privacy: BECS resolves the inherent tension between market transparency and user privacy by architecturally decoupling the public visibility of resource status from the identities of participants. Global transparency of resource availability is an essential requirement for any decentralized resource marketplace, as it enables requesters to efficiently discover potential providers. BECS ensures that such transparency does not compromise user privacy by incorporating a robust pseudonym scheme, where all transactions are performed through temporary, unlinkable identifiers. Consequently, it prevents both external and internal adversaries from constructing long-term behavioral profiles or linking a user’s activities across multiple transactions, thereby achieving an effective balance between market efficiency and strong privacy preservation.
VI-C Pseudonym and Certificate Security
The pseudonym scheme is equipped with specific cryptographic properties to mitigate the capabilities of both adversaries, $\mathcal{A}$ and $\mathcal{B}$, across a range of attack scenarios.
1) Uniqueness and Authenticity: This property mitigates impersonation attacks launched by both adversaries, $\mathcal{A}$ and $\mathcal{B}$. The zero-knowledge protocol-based generation process ensures that only the legitimate owner of a private key $d$ can generate a valid pseudonym and its corresponding proof. As a result, it becomes computationally infeasible for an adversary to forge a valid pseudonym for another user, thereby ensuring that each identity within a transaction is authentic.
2) Unforgeability: Certificate unforgeability directly prevents both adversaries from obtaining illegitimate trust within the system. Since generating a valid certificate requires access to the private key $b$ of a 6GBS, no adversary can fabricate a fraudulent certificate, thereby preventing impersonation of certified and trustworthy nodes.
3) Unlinkability: This property plays a critical role in defending against the privacy-compromising objectives of both adversary types.
  • Against the external adversary $\mathcal{A}$, the use of fresh, random nonces $\gamma$ in each session ensures that pseudonyms appear as independent and uncorrelated values on the public channel. This renders traffic analysis and activity inference through eavesdropping ineffective.
  • Against the malicious internal node $\mathcal{B}$, the zero-knowledge-based interactions and session-specific pseudonyms prevent it from linking its current transaction partner to any previous or future transactions. This prevents malicious peers from constructing a user’s long-term behavioral profile.
4) Traceability: This feature is specifically designed as a deterrent and mitigation strategy against the malicious internal node $\mathcal{B}$. While unlinkability safeguards users from peer inference, authorized traceability ensures that when malicious behavior is verified, the 6GBSs can revoke the anonymity of the offending node and expel it from the system. This accountability mechanism deters internal misbehavior without compromising user anonymity.
VI-D Privacy-Performance Trade-off
The inclusion of privacy entropy as a simultaneous optimization objective introduces a fundamental trade-off with performance-oriented metrics. A strategy optimized exclusively for performance, such as minimizing latency or energy consumption, would inevitably favor deterministic offloading decisions and thus produce predictable, low-entropy patterns. As established in the threat model, these predictable patterns constitute the primary attack surface exploited by $\mathcal{C}$ to compromise user privacy. Maximizing privacy entropy necessitates introducing a degree of non-determinism into the allocation process to obscure these patterns, thereby obfuscating the correlation between offloading decisions and the user’s state.
This inherent conflict is not a limitation to be eliminated but rather a fundamental challenge that must be intelligently managed, validating the necessity of a MOOP. The NSGA-III-KDR algorithm is designed not to find a single utopian solution but to identify the Pareto-optimal front, which consists of a set of non-dominated solutions, each representing a quantitatively optimal and practically meaningful trade-off between robust privacy and high performance. The final policy is then selected using TOPSIS to align with specific application requirements, thereby transforming the trade-off into a mechanism for flexible and context-aware optimization.
VI-E Computational Complexity of NSGA-III-KDR
@@
In BECS, NSGA-III-KDR uses Non-KDR-dominated sort instead of Non-Dominated Sort from NSGA-III to enhance solution diversity, while maintaining the same computational complexity. Namely, with $L$ optimization objectives and a population size of $X$, NSGA-III-KDR exhibits a computational complexity of $\mathcal{O}(LX^{2})$. Specifically, calculating the kernel distance for each solution and determining the angle between any two solutions each incur a complexity of $\mathcal{O}(L)$. Each solution’s distance to the ideal point is calculated individually, contributing to a complexity of $\mathcal{O}(LX)$. For dominance judgment, each pair of solutions undergoes one angle calculation and comparison, leading to an overall complexity of $\mathcal{O}(LX^{2})$. Thus, the computational complexity of each iteration can be approximated as $\mathcal{O}(LX^{2})$. Compared to the NSGA-III-based scheme [6], NSGA-III-KDR offers more diverse solutions, improving the probability of superior computing allocation strategies.
@@ VII Simulation Results and Analysis
TABLE III: Key Parameters
@@
[…]
@@
In this section, we initially compare the performance of the proposed NSGA-III-KDR with NSGA-III and MOEA/D applied in computing allocation, as well as the NSGA-II-SDR that inspired us. Subsequently, we test the performance of four algorithms in optimizing the proposed computing allocation MOOP. Finally, we simulate the performance of the proposed pseudonym scheme. The configurations of critical parameters are detailed in Table III. The code will be available at https://github.com/QuinYim/BECS.
@@ VII-A Simulation of Proposed NSGA-III-KDR
In this part, we compare the proposed NSGA-III-KDR with state-of-the-art evolutionary algorithms NSGA-III [6], MOEA/D [7], and NSGA-II-SDR [27], utilizing the PlatEMO platform [40]. The widely used SDTLZ [25], MaF [42], and SMOP [43] test suites are employed as the benchmarks. Meanwhile, IGD [44] and PD [45] are selected as performance evaluation metrics. IGD comprehensively measures the convergence and diversity of the algorithms, where a smaller IGD value indicates better performance. PD primarily reflects the diversity of the algorithms, with a larger PD value indicating greater population diversity. The crossover probability is set to 1, the mutation probability is set to 1/D, and their distribution indicator is set to 20, where D represents the length of the decision variable. All algorithms are executed for 1000 iterations over 30 independent times on different test problems, and the average values are taken. The performance of four algorithms is compared under different numbers of objectives and benchmarks, as shown in Table IV, where “++”, “-”, and “==” indicate that the result is significantly better, significantly worse, and statistically similar to that obtained by NSGA-III-KDR, respectively.
TABLE IV: IGD and PD value of NSGA-III, MOEA/D, NSGA-II-SDR, and NSGA-III-KDR on SDTLZ1, SDTLZ2, MaF1, MaF2, SMOP1, and SMOP2 with 5, 10, and 15 objectives. The best result in each row is highlighted.
@@
It can be concluded from the experimental results that NSGA-III-KDR has the strongest competitiveness, achieving the best IGD and PD numbers of 10 and 15, respectively. This demonstrates that NSGA-III-KDR achieves better performance compared to the other two algorithms, especially in terms of population diversity, offering a richer set of solutions for computing allocation.